OpenStudy (anonymous):

Find the equation of the line that passes through (1, -1) and (10, 2).

OpenStudy (luigi0210):

First what you could do is find the slope

OpenStudy (anonymous):

idk how to do that im really bad with math could you help me get the answer?

OpenStudy (luigi0210):

Well you would get the slope by using y2-y1/x2-x1where y2=2, y1=-1, x2=10 and x1=1

OpenStudy (anonymous):

slope is 4

OpenStudy (luigi0210):

\[\frac{ (2)-(-1) }{ (10)-(1) }=4?\]

OpenStudy (anonymous):

x1 is 1 x2 is -1 y1 is 10 y2 is 2 correct?

OpenStudy (anonymous):

could you just give me the answer plz?

OpenStudy (luigi0210):

Well the slope is 1/3.. and all you do is plug this into the point slope form formula. so Using the point (1, -1) we get \[y+1=\frac{ 1 }{ 3 }(x-1)\] Should I continue or can you finish it from here?

OpenStudy (anonymous):

continue plz

OpenStudy (luigi0210):

Okay,now distribute the factors out: